OpenStudy (anonymous):

Write the equation of the parabola in vertex form vertex (2,1), point (3,-4)

OpenStudy (anonymous):

y=a(x-2)^2 +1 -4=a(3-2)^2+1 -5=a y=-5(x-2)^2 + 1